Q5. (3 marks) Solve the initial value problem (x + 3)y + y = (x + 3)3, y(0) = 0, and sketch the solution
Answer #1

Given equation is:

(x+3)y' +y = (x+3)^3

We can write it as:

y+ = (z + 3) (z+3)

Now this is a linear differential equation of first order of the form

y' +P(x)y = Q(x)

In this P(x)=\frac{1}{x+3} , Q(x)= (x+3)^2

Integrating factor of this equation is:

I.F = e^{\int P(x)dx}

= e^{\int \frac{1}{(x+3)}dx}

= e^{ln(x+3)} = (x+3)

Solution of this differential equation is:

y*(I.F) = \int Q(x)*(I.F)dx+ C

y(x+3) = \int (x+3)^2*(x+3)dx+ C

y(x+3) = \int (x+3)^3dx+ C

y(x+3) = \frac{(x+3)^4}{4}+ C

We can write it as:

(x+3)y = \frac{(x+3)^4}{4}+ C

As y(0) = 0

(0+3)*0 = \frac{(0+3)^4}{4}+ C

0 = \frac{81}{4}+ C

So C =- \frac{81}{4}

SO solution of given equation is:

(x+3)y = \frac{(x+3)^4}{4}-\frac{81}{4}